Meganla:

The ratio of the number of cups of red paint to the number of cups of yellow paint that Garcia mixed to make a shade of orange paint is 4:1. Part A: Write a sentence to describe the number of cups of red paint that Garcia mixed with each cup of yellow paint. (4 points) Part B: How many cups of yellow paint should Garcia add to each cup of red paint to make the shade of orange? Explain your reasoning. (4 points) Part C: If Garcia purchased 6 cups of red paint for $42, what is the unit price per cup of red paint? (2 points)

Meganla:

Part A: Garcia mixed 4 cups of RED with each cup of YELLOW. Part B: Garcia should add 1/4 cup of YELLOW paint to each cup of RED paint. 4:4=1 1:4=1/4 1:1/4 Part C: 6 cups of RED paint = 42 1 cups of RED paint = 7 42 : 6 = 7

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Meganla Part A: Garcia mixed 4 cups of RED with each cup of YELLOW. Part B: Garcia should add 1/4 cup of YELLOW paint to each cup of RED paint. 4:4=1 1:4=1/4 1:1/4 Part C: 6 cups of RED paint = 42 1 cups of RED paint = 7 42 : 6 = 7 \(\color{#0cbb34}{\text{End of Quote}}\) U got everything right but,But i think u should go more into detail with (A) like perhaps- For every one cup of yellow paint, Garcia added four cups of red paint, which is a ratio of 4:1.

Meganla:

Thank you

Hoodmemes:

np